calculate the molar mass of magnesium sulfate
1
Expert's answer
2017-03-10T10:34:05-0500
M(MgSO4) = Ar(Mg) + Ar(S) + 4Ar(O) = 24 + 32 + 4·16 = 120 g/mol.

Answer:
120 g/mol